OpenStudy (idealist10):

How to simplify e^(x+(2/3)x^3)?

OpenStudy (idealist10):

@Nurali @dan815 @ganeshie8

OpenStudy (idealist10):

@wio @jhonyy9 @nincompoop @uri

OpenStudy (anonymous):

it looks simplified to me.

OpenStudy (anonymous):

\[ e^{x+\frac 23 x^3} \]You can't simplify the exponent.

OpenStudy (idealist10):

Okay, then. It's because I want to solve the problem about finding the general solution of xy'+(1+2x^2)y=(x^3)(e^(-x^2)) so I figured out that the integrating factor is e^(x+(2/3)x^3).

OpenStudy (idealist10):

So do I multiply that by the whole differential equation since you said it can't be simplified more?

OpenStudy (anonymous):

yes

OpenStudy (idealist10):

Wait a minute. Please don't leave. Let me work it out then.

OpenStudy (idealist10):

How do I integrate (x^3)(e^(x+(2/3)x^3-x^2))?

OpenStudy (anonymous):

I don't think it has an elementary anti-derivative.

OpenStudy (idealist10):

Am I doing something wrong then?

OpenStudy (idealist10):

@Abhisar @johnweldon1993 @Preetha @paki

OpenStudy (idealist10):

@charlotte123 @dan815

OpenStudy (anonymous):

don't you need to divide by \(x\) first?

OpenStudy (anonymous):

I thought that \(xy'\) is wrong since you want just \(y'\) with no coefficient.

OpenStudy (idealist10):

How?

OpenStudy (anonymous):

\[ y' +(x^{-1}+2x)y = x^2e^{-x^2} \]

OpenStudy (anonymous):

That would make your integrating factor: \[ \exp\left(\int x^{-1}+2x\;dx\right) = \exp(\ln(x)+x^2) = e^{\ln x+x^2} = xe^{x^2} \]

OpenStudy (johnweldon1993):

So lets see... \[\large xy' + (1 + 2x^2)y = (x^3)(e^{-x^2})\] ***divide everything by x to make y' by itself \[\large \frac{dy(x)}{dx} + (\frac{1}{x} + 2x)y(x) = e^{-x^2} x^2\] And from there it looks like the intergrating factor should be \(\large e^{\int(\frac{1}{x} + 2x)dx}\)

OpenStudy (idealist10):

Let me try that, wio. Please don't leave.

OpenStudy (johnweldon1993):

yes perfect as wio has it!

OpenStudy (anonymous):

The constant of integration would make it \(Cxe^{x^2}\).

OpenStudy (idealist10):

Yes, I got it! The integrating factor is xe^(x^2) like wio said. I've got the right answer. Thank you so much, wio! Since you told me to divide everything by x at first, it made everything so much easier!

OpenStudy (anonymous):

You always need to get the \(y'\) term to have a coefficient of 1.

OpenStudy (anonymous):

It's standard form.

OpenStudy (idealist10):

I see.